# ToDo (#1): #2

### v2

@@
Abstract: This Preliminary Design Report (PDR) describes the IsoDAR electron-antineutrino source. Volumes I and II are site-independent and describe the cyclotron driver providing a 10 mA proton beam, and the medium energy beam transport line and target, respectively. Volume III describes the installation at the Yemilab underground laboratory in South Korea. The IsoDAR driver and target will produce a mole of electron-antineutrinos over the course of five years. Paired with a kton-scale liquid scintillator detector, it will enable an impressive particle physics program including searches for new symmetries, new interactions and new particles. Here in Volume I, we describe the driver, which includes the ion source, low energy beam transport, and cyclotron. The latter features radiofrequency quadrupole (RFQ) direct axial injection and represents the first accelerator purpose-built to make use of vortex motion.
[…]
@@
${}^{1}$on Beam Applications, 3 chemin du cyclotron Louvain-la-Neuve, BELGIUM
${}^{2}$assachusetts Institute of Technology, 77 Massachusetts Ave. Cambridge, MA, USA
@@
Figure 1.1: Schematic of the IsoDAR experiment installed at the Yemilab facility in South Korea. From right to left: The compact cyclotron with RFQ injector, the medium energy beam transport, the target, and the LSC neutrino detector. From [1].
@@
IsoDAR will be the first high-intensity, proton-driven source to be located alongside an underground multi-kiloton-scale scintillator detector. The first version of IsoDAR will be constructed at the Yemilab facility [2] within the Handuk mine, located in South Korea where the “Liquid Scintillator Counter” (LSC) is planned to be installed [3]. The novel source, consisting of a cyclotron, beam transport line, and a target/sleeve is shown in Fig. 1.1. In five calendar years (four live years), the source will produce 1.67 million $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ (inverse beta decay or ”IBD”) events and 7000 $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$ (elastic scattering or ES) events in the LSC that can be used for searches for new physics. More than a mole of neutrons and monoenergetic photons are also produced in the target, and these may act as progenitors for new particles that enter the LSC and leave signatures from interactions and decays. IsoDAR can address a broad range of new physics manifested as new symmetries, new interactions and new particles that is inaccessible to traditional experiments. For further discussion see Refs. [4, 5].
@@
This document is Volume 1 in a three-volume Preliminary Design Report (PDR) for IsoDAR that provides the engineering solutions necessary to establish credible solutions to the key issues and cost-drivers identified during the past decade of conceptual design and R&D. The contents of the PDR will cover the Technical Facility: the cyclotron, the beam transport and the target/sleeve. This document builds on the structure and text of the Conceptual Design Report (CDR) of the IsoDAR Technical Facility [6]. This step of producing a PDR substantially reduces risk related to IsoDAR approval: the remaining detailed engineering that will follow will build on these results and be straightforward to estimate in scope and cost.
@@
Volume 1 describes the IsoDAR proton driver. This consists of a “front end” that delivers ions to a cyclotron for acceleration and extraction and the cyclotron itself. This document presupposes a basic understanding of cyclotrons; for an introduction to cyclotrons and a cost/benefit analysis of this choice of machine for IsoDAR, see the discussion in the CDR.
@@
The 60 MeV/amu IsoDAR Proton Driver is exceptional in that it is designed and engineered to produce an order of magnitude higher intensity proton beam at extraction than commercial machines of similar energy. Achieving this high power necessitated three key breakthroughs. First, the cyclotron accelerates $\mathrm{H}_{2}^{+}$, mitigating “space charge effects” (electromagnetic interactions between beam particles). Second, the beam is axially injected through an RFQ, acting as a high-efficiency buncher. Third, the design harnesses vortex motion, a complex effect previously observed at PSI Injector II [7, 8], inducing stable spiraling motion in the high-current accelerated beam, significantly reducing beam growth. For more concerning the accelerator physics that shapes the IsoDAR design, see Ref. [9, 10].
@@
Chapter 1:
(This chapter) proceeds with an overview of the physics searches that will be possible with IsoDAR, followed by an overview of the requirements for the proton driver.
Chapter 2:
Describes engineering related to the IsoDAR front end.
Chapter 3:
Describes engineering related to the general IsoDAR cyclotron design.
Chapter 4:
Describes specifics of the design related to installation underground.
@@ 1.1 The IsoDAR Physics Case
Figure 1.2: Energy dependence of fluxes. Left: $\bar{\nu}_{e}$, unit normalized; Middle: photons per proton on target (POT) from excited nuclei. Right: kinetic energy of neutrons at production in target and sleeve, normalized per proton.
IsoDAR offers a rich BSM physics program if paired with a kton-scale liquid scintillator detector like the proposed LSC at Yemilab [3]. The exceptional reach is owed to the high flux rates of $\bar{\nu}_{e}$, photons, and neutrons from the $\approx 2\cdot 10^{24}$ protons on target over 5 years (cf. Fig. 1.2), and placing the source very close to the detector. Here we briefly discuss the highlights of the IsoDAR BSM physics program enabled by a 2.26 kton detector like the LSC in three categories: “New Particles” - experimentally and theoretically motivated; “New Symmetries” - CP-violation and the QCD Axion; And “New Interactions” - Precision Electroweak Physics. See Refs. [4, 11, 5] for further details.
@@ 1.1.1 Fluxes in IsoDAR
The $\bar{\nu}_{e}$ flux stems from a single beta-decay of ${}^{8}$i, which is produced by neutron capture on ${}^{7}$i and is well-understood. It peaks at 6 MeV, which puts it above backgrounds from the environment. These are clear advantages over reactor-neutrino fluxes. Interactions in the detector at these neutrino energies are exclusively inverse-beta-decay (IBD) and elastic scattering (ES). Cross sections for these reactions are known with uncertainties at $<1$% at these energies.
Neutrons in the IsoDAR target are primarily produced by spallation interactions of the beam protons with the beryllium target hemispheres. Additional secondary neutrons stem from interactions of the primary neutrons with beryllium in the target and sleeve. We model the processes in the target with Geant4 using an accurate representation of the latest target and sleeve design (see Refs. [12]). We show the resulting neutron energy spectrum in Fig 1.2 (right). Most neutrons capture on the ${}^{7}$i in the sleeve. To ensure $<5$ neutrons per year enter the LSC veto region, the target is further shielded using boron-loaded concrete and steel [13].
The third flux we consider is that of photons as seen in Fig.1.2 (middle). Here, one can observe a line structure that stems from the transitions of excited nuclei and and underlying smooth continuum due to bremsstrahlung. Again, the spectrum was generated using Geant4 using the full IsoDAR Target geometry [5].
@@ 1.1.2 New Particles: Experimentally and Theoretically Motivated
Figure 1.4: Left: Example $X\rightarrow\nu\bar{\nu}$ signatures for 0.5% photon mixing with ${}^{8}$i-produced $\bar{\nu}_{e}$ background (gray). Right: $X$ mass and $X$-to-photon production ratio 90% CL sensitivity [4].
Another example of IsoDAR’s capability to look for new particles is the “$X$,” which is a low mass mediator that would couple to the photons produced in the target [4]. The $X$ would be a light new particle that decays to $\bar{\nu}\nu$ and would leave a distinct signature above background in the IBD spectrum. Fig.1.4 (left) displays resulting signatures for various masses (see color bar) assuming 0.5% $X$-to-photon production. The IBD background from $\bar{\nu}_{e}$ stemming from ${}^{8}$i decays (gray) constitutes the primary background and has a distinct shape. The sensitivity is depicted in Fig.1.4 (right), where the blue region is excluded in the case of the SM. No existing limits are shown because this search is first-of-its-kind. A 17 MeV $X$ particle is consistent with the Atomki anomaly.
@@ 1.1.3 New Symmetries: Strong CP Violation and the QCD Axion
Figure 1.5: IsoDAR sensitivity for axions (a) with mass $m_{a}$ coupling ($g$) to electrons (lines) [5]. Gold dots–axion allowed; Teal–excluded regions; Gray–parameters with astrophysical implications.
The neutron beam/bottle experiment lifetime discrepancy [17, 18, 19, 20] motivates neutron–dark sector couplings. Through $n\rightarrow n^{\prime}\rightarrow n$ transitions, we may see neutrons disappearing from the target and reappear in the LSC. The underground environment provides a low background and the LSC’s size further adds to the sensitivity [11]. Alternatively, if the $n^{\prime}$ are Majorana, then $n\rightarrow n^{\prime}\rightarrow\bar{n}$ may occur, depositing $\sim$2 GeV of energy in the LSC upon annihilation. This is a $B$ violating effect with implications for the early universe [21].
The photons and electrons in the target may also couple to axions and IsoDAR can probe uncharted territory with astrophysical and cosmological implications [5]. In Fig. 1.5 we show QCD axions coupling to electrons.
@@ 1.1.4 New Interactions: Tests via Precision Electroweak Physics
IsoDAR will record about 7000 elastic scatter events over its 4-year accumulated up-time. This constitutes the world’s largest sample at low momentum transfer. LEP and LHC data predict rate and energy dependency very precisely within the standard model [22]. With the normalization determined by the $\approx 1.7$M IBD events in the LSC, we will be able to measure $\sin^{2}\theta_{W}$ with a precision of 1.9% (see Fig.1.6 left).
@@
Figure 1.6: Left: Weak Mixing Angle expectation for IsoDAR compared to present results and the future DUNE expectation. Right: expectation for IsoDAR in NSI parameter space. See Ref. [4].
The accurately predicted energy dependence of the ES sample enables a precise phenomenological search for non-standard neutrino interactions (NSIs). This powerful search for lepton-lepton couplings complements the lepton-quark coupling searches by the active, global coherent neutrino scattering program. The SM right- and left-handed couplings, $g_{R}$ and $g_{L}$, are altered by NSI corrections for electron flavor antineutrino interactions with electrons (hence ”ee”) to $\tilde{g}_{R}=g_{R}+\varepsilon_{ee}^{eR}$ and $\tilde{g}_{L}=g_{L}+\varepsilon_{ee}^{eL}$. The SM expectation for IsoDAR@Yemilab in the parameter space of these deviations is seen in Fig. 1.6 (right), where red assumes background reduction using Cherenkov light, and green assumes only scintillation light reconstruction. This can be compared to the world result [23] in black, which is dominated by $\nu_{e}$–electron ES scattering from beam dump experiments and, hence, is rotated with respect to IsoDAR, which has a pure $\bar{\nu}_{e}$ flux.
@@ 1.2 Requirements for the Proton Driver
As shown in Fig. 1.1, the Proton Driver consists of a 60 MeV/amu cyclotron accelerating and extracting 5 electrical milliamperes of $\mathrm{H}_{2}^{+}$, a transport line that strips the $\mathrm{H}_{2}^{+}$ ions close to the extraction point, analyzes the stripped beam to monitor the health of the stripping foil, and transports the 10 milliamperes of protons to the target. The target consists of three nested hemispherical beryllium shells, the outermost being 20 cm in diameter, cooled with heavy water flowing around the shell structure. The proton beam is shaped, with quadrupoles and wobbler magnets to produce a distribution that allows for optimal heat transfer to the cooling water. Neutrons produced in the target system are moderated and flood into a roughly-spherical sleeve, about 1 meter in diameter, containing a mixture of beryllium chips and highly-enriched ${}^{7}$i. The beta-decay of the ${}^{8}$i resulting from the capture of neutrons produces the neutrinos of interest in the experiment.
The principal requirements for the accelerator system are to reliably produce a 10 mA beam of protons in the most efficient manner possible with the lowest possible loss of particles in the acceleration process. Experience from the cyclotron system at PSI [24] are that keeping the total power of beam loss inside the cyclotron vault to less than 200 watts still allows access for component maintenance. Considering that the total beam power is 600 kilowatts, one sees that this is a very stringent requirement.
As mentioned above, Volume 1 of this PDR addresses the physics and engineering aspects of the 60 MeV/amu cyclotron. The final chapter addresses the complex issue of transport and assembly of the large and heavy cyclotron components in the underground environment at Yemilab.
Subsequent volumes will address the details of the beam transport system and the target, along with shielding requirements for the environment of the target including the prevention of fast neutrons from reaching the fiducial volume of the liquid scintillator detector and the underground installation procedure, power and cooling, HVAC, and other site-specific topics.

### v3

@@
Abstract: This Preliminary Design Report (PDR) describes the IsoDAR electron-antineutrino source in two volumes which are mostly site-independent and describe the cyclotron driver, providing a 10 mA proton beam, and the medium energy beam transport line and target, respectively. The IsoDAR driver and target will produce a mole of electron-antineutrinos over the course of five years. Paired with a kton-scale liquid scintillator detector, it will enable an impressive particle physics program including searches for new symmetries, new interactions and new particles. Here in Volume I, we describe the driver, which includes the ion source, low energy beam transport, and cyclotron. The latter features radiofrequency quadrupole (RFQ) direct axial injection and represents the first accelerator purpose-built to make use of so-called vortex motion.
Section Editors (in alphabetical order): M. Abs1, J.R Alonso2, J.M. Conrad2, S.J.B. Engebretson2, E. Forton1, A. Herrod1, D. Joassin1, J. Moon2, S. de Neuter1, E. Van der Kraaij1, G. Wéry1, D. Winklehner2, E. Winkler2
@@
1Ion Beam Applications, 3 chemin du cyclotron Louvain-la-Neuve, BELGIUM
2Massachusetts Institute of Technology, 77 Massachusetts Ave. Cambridge, MA, USA
@@
Figure 1.1: Schematic of the IsoDAR experiment installed at the Yemilab facility in South Korea. From right to left: The compact cyclotron with RFQ injector, the medium energy beam transport, the target, and the LSC neutrino detector. From Ref. [calvo_analyzing_2023].
@@
IsoDAR will be the first high-intensity, proton-driven source to be located alongside an underground multi-kiloton-scale scintillator detector. The first version of IsoDAR will be constructed at the Yemilab facility [alonso_isodaryemilab_2022] within the Handuk mine, located in South Korea where the “Liquid Scintillator Counter” (LSC) [^1] The LSC was renamed $\nu$Eye in March 2025, but for consistency throughout this document, we will refer to the detector as the LSC. is planned to be installed [seo_physics_2023]. The novel source, consisting of a cyclotron, beam transport line, and a target/sleeve is shown in Fig. 1.1. In five calendar years (four live years), the source will produce 1.67 million $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ (inverse beta decay or ”IBD”) events and 7000 $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$ (elastic scattering or ES) events in the LSC that can be used for searches for new physics. More than a mole of neutrons and mono-energetic photons are also produced in the target, and these may act as progenitors for new particles that enter the LSC and leave signatures from interactions and decays. IsoDAR can address a broad range of new physics manifested as new symmetries, new interactions and new particles that is inaccessible to traditional experiments. For further discussion see Refs. [alonso_neutrino_2022, waites_axionlike_2023].
@@
This document is Volume I in a two-volume Preliminary Design Report (PDR) for IsoDAR that provides the engineering solutions necessary to establish credible solutions to the key issues and cost-drivers identified during the past decade of conceptual design and R&D. The contents of the PDR will cover the Technical Facility: the cyclotron, the beam transport and the target/sleeve. This document builds on the structure and text of the Conceptual Design Report (CDR) of the IsoDAR Technical Facility [abs_isodarkamland_2015]. This step of producing a PDR substantially reduces risk related to IsoDAR approval: the remaining detailed engineering that will follow will build on these results and be straightforward to estimate in scope and cost.
@@
Volume I describes the IsoDAR proton driver. This consists of a “front end” that delivers ions to a cyclotron for acceleration and extraction and the cyclotron itself. This document presupposes a basic understanding of cyclotrons; for an introduction to cyclotrons and a cost/benefit analysis of this choice of machine for IsoDAR, see the discussion in the CDR.
@@
The 60 MeV/amu IsoDAR Proton Driver is exceptional in that it is designed and engineered to produce an order of magnitude higher current proton beams at extraction than commercial machines of similar energy. Achieving this high power necessitated three key breakthroughs. First, the cyclotron accelerates $\mathrm{H}_{2}^{+}$, mitigating “space charge effects” (electromagnetic interactions between beam particles). Second, the beam is axially injected through an RFQ, acting as a high-efficiency buncher. Third, the design harnesses vortex motion, a complex effect previously observed at PSI Injector II [stetson:vortex, baumgarten:vortex1], inducing stable spiraling motion in the high-current accelerated beam, significantly reducing beam growth. For more concerning the accelerator physics that shapes the IsoDAR design, see Ref. [winklehner_order--magnitude_2022, winklehner_high_2018].
@@
  • Chapter 1: (This chapter) proceeds with an overview of the physics searches that will be possible with IsoDAR, followed by an overview of the requirements for the proton driver.
  • Chapter 2: Describes engineering related to the IsoDAR front end.
  • Chapter 3: Describes engineering related to the general IsoDAR cyclotron design.
  • Chapter 4: Describes specifics of the design related to installation underground.
@@ 1.1 The IsoDAR Physics Case
Figure 1.2: Energy dependence of fluxes. Left: $\bar{\nu}_{e}$, unit normalized; Middle: photons per proton on target (POT) from excited nuclei [alonso_neutrino_2022]. Right: kinetic energy of neutrons at production in target and sleeve, normalized per proton.
IsoDAR offers a rich BSM physics program if paired with a kton-scale liquid scintillator detector like the proposed LSC at Yemilab [seo_physics_2023]. The exceptional reach is owed to the high flux rates of $\bar{\nu}_{e}$, photons, and neutrons from the $\approx 2\cdot 10^{24}$ protons on target over 5 years (cf. Fig. 1.2), and placing the source very close to the detector. Here we briefly discuss the highlights of the IsoDAR BSM physics program enabled by a 2.26 kton detector like the LSC in three categories: “New Particles” - experimentally and theoretically motivated; “New Symmetries” - CP-violation and the QCD Axion; And “New Interactions” - Precision Electroweak Physics. See Refs. [alonso_neutrino_2022, Hostert:2022ntu, waites_axionlike_2023] for further details.
@@ 1.1.1 Fluxes in IsoDAR
The $\bar{\nu}_{e}$ flux stems from a single beta-decay of 8Li, which is produced by neutron capture on 7Li and is well-understood. It peaks at 6 MeV, which puts the majority of the events above backgrounds from the environment. These are clear advantages over reactor-neutrino fluxes. Interactions in the detector at these neutrino energies are exclusively inverse-beta-decay (IBD) and elastic scattering (ES). Cross sections for these reactions are known with uncertainties at $<1$% at these energies.
Neutrons in the IsoDAR target are primarily produced by spallation interactions of the beam protons with the beryllium target hemispheres. Additional secondary neutrons stem from interactions of the primary neutrons with beryllium in the target and sleeve. We model the processes in the target with Geant4 using an accurate representation of the latest target and sleeve design (see Refs. [bungau_optimizing_2019]). We show the resulting neutron energy spectrum in Fig 1.2 (right). Most neutrons capture on the 7Li in the sleeve. To ensure $<5$ neutrons per year enter the LSC veto region, the target is further shielded using boron-loaded concrete and steel [bungau_shielding_2020].
The third flux we consider is that of photons as seen in Fig.1.2 (middle). Here, one can observe a line structure that stems from the transitions of excited nuclei and and underlying smooth continuum due to bremsstrahlung. Again, the spectrum was generated using Geant4 using the full IsoDAR Target geometry [waites_axionlike_2023].
@@ 1.1.2 New Particles: Experimentally and Theoretically Motivated
Figure 1.4: Left: Example $X\rightarrow\nu\bar{\nu}$ signatures for 0.5% photon mixing with 8Li-produced $\bar{\nu}_{e}$ background (gray). Right: $X$ mass and $X$-to-photon production ratio 90% CL sensitivity. From Ref. [alonso_neutrino_2022].
Another example of IsoDAR’s capability to look for new particles is the “$X$,” which is a low mass mediator that would couple to the photons produced in the target [alonso_neutrino_2022]. The $X$ would be a light new particle that decays to $\bar{\nu}\nu$ and would leave a distinct signature above background in the IBD spectrum. Fig.1.4 (left) displays resulting signatures for various masses (see color bar) assuming 0.5% $X$-to-photon production. The IBD background from $\bar{\nu}_{e}$ stemming from 8Li decays (gray) constitutes the primary background and has a distinct shape. The sensitivity is depicted in Fig.1.4 (right), where the blue region is excluded in the case of the SM. No existing limits are shown because this search is first-of-its-kind. A 17 MeV $X$ particle is consistent with the Atomki anomaly.
@@ 1.1.3 New Symmetries: Strong CP Violation and the QCD Axion
Figure 1.5: IsoDAR sensitivity for axions (a) with mass $m_{a}$ coupling ($g$) to electrons (lines) [waites_axionlike_2023]. Gold dots–axion allowed; Teal–excluded regions; Gray–parameters with astrophysical implications.
The neutron beam/bottle experiment lifetime discrepancy [Nico:2004ie, Yue:2013qrc, Serebrov:2017bzo, UCNt:2021pcg] motivates neutron–dark sector couplings. Through $n\rightarrow n^{\prime}\rightarrow n$ transitions, we may see neutrons disappearing from the target and reappear in the LSC. The underground environment provides a low background and the LSC’s size further adds to the sensitivity [Hostert:2022ntu]. Alternatively, if the $n^{\prime}$ are Majorana, then $n\rightarrow n^{\prime}\rightarrow\bar{n}$ may occur, depositing $\sim$2 GeV of energy in the LSC upon annihilation. This is a $B$ violating effect with implications for the early universe [Berezhiani:2020vbe].
The photons and electrons in the target may also couple to axions and IsoDAR can probe uncharted territory with astrophysical and cosmological implications [waites_axionlike_2023]. In Fig. 1.5 we show QCD axions coupling to electrons.
@@ 1.1.4 New Interactions: Tests via Precision Electroweak Physics
IsoDAR will record about 7000 elastic scatter events over its 4-year accumulated up-time. This constitutes the world’s largest sample at low momentum transfer. LEP and LHC data predict rate and energy dependency very precisely within the standard model [ParticleDataGroup:2020ssz]. With the normalization determined by the $\approx 1.7$M IBD events in the LSC, we will be able to measure $\sin^{2}\theta_{W}$ with a precision of 1.9% (see Fig.1.6 left).
@@
Figure 1.6: Left: Weak Mixing Angle expectation for IsoDAR compared to present results and the future DUNE expectation. Right: expectation for IsoDAR in NSI parameter space. From Ref. [alonso_neutrino_2022].
The accurately predicted energy dependence of the ES sample enables a precise phenomenological search for non-standard neutrino interactions (NSIs). This powerful search for lepton-lepton couplings complements the lepton-quark coupling searches by the active, global coherent neutrino scattering program. The SM right- and left-handed couplings, $g_{R}$ and $g_{L}$, are altered by NSI corrections for electron flavor antineutrino interactions with electrons (hence ”ee”) to $\tilde{g}_{R}=g_{R}+\varepsilon_{ee}^{eR}$ and $\tilde{g}_{L}=g_{L}+\varepsilon_{ee}^{eL}$. The SM expectation for IsoDAR@Yemilab in the parameter space of these deviations is seen in Fig. 1.6 (right), where red assumes background reduction using Cherenkov light, and green assumes only scintillation light reconstruction. This can be compared to the world result [guzzo] in black, which is dominated by $\nu_{e}$–electron ES scattering from beam dump experiments and, hence, is rotated with respect to IsoDAR, which has a pure $\bar{\nu}_{e}$ flux.
@@ 1.2 Requirements for the Proton Driver
As shown in Fig. 1.1, the Proton Driver consists of a 60 MeV/amu cyclotron accelerating and extracting 5 electrical milliamperes of $\mathrm{H}_{2}^{+}$, a transport line that strips the $\mathrm{H}_{2}^{+}$ ions close to the extraction point, analyzes the stripped beam to monitor the health of the stripping foil, and transports the 10 milliamperes of protons to the target. The target consists of three nested hemispherical beryllium shells, the outermost being 20 cm in diameter, cooled with heavy water flowing around the shell structure. The proton beam is shaped, with quadrupoles and wobbler magnets to produce a distribution that allows for optimal heat transfer to the cooling water. Neutrons produced in the target system are moderated and flood into a roughly-spherical sleeve, about 1 meter in diameter, containing a mixture of beryllium chips and highly-enriched 7Li. The beta-decay of the 8Li resulting from the capture of neutrons produces the neutrinos of interest in the experiment.
The principal requirements for the accelerator system are to reliably produce a 10 mA beam of protons in the most efficient manner possible with the lowest possible loss of particles in the acceleration process. Particularly above the Coulomb threshold. The experience of the cyclotron system at PSI [Stammbach] is that keeping the total power of beam loss inside the cyclotron vault to less than 200 W still allows access for component maintenance. Considering that the total beam power is 600 kW, one sees that this is a very stringent requirement.
As mentioned above, Volume I of this PDR addresses the physics and engineering aspects of the 60 MeV/amu cyclotron. The final chapter addresses the complex issue of transport and assembly of large and heavy cyclotron components in the underground environment at Yemilab.
A forthcoming Volume II will address the details of the beam transport system and the target, along with shielding requirements for the environment of the target, including the prevention of fast neutrons from reaching the fiducial volume of the liquid scintillator detector.
1.3 Overview of the Accelerator Layout
Figure 1.7: Overview of the components of the HCHC-60 Cyclotron System. Only one RF cavity is shown to not clutter the image.
To guide the reader through this PDR and, indeed, the path of an ion from genesis in the ion source to the full 60 MeV/amu and extraction, we provide an overview in Fig. 1.7. The main components of the system, which will be described in much detail in the following chapters, are:
  • The ion source, providing 10 mA of $\mathrm{H}_{2}^{+}$ at an emittance $<0.1\pi$-mm-mrad.
  • The Low Energy Beam Transport (LEBT), shaping and guiding the beam to the RFQ and providing diagnostics, and chopping for commissioning and personnel- and machine protection.
  • The Radiofrequency Quadrupole (RFQ), embedded in the cyclotron yoke, to bunch and pre-accelerate the beam to 35 keV/amu.
  • The central region with electrostatic quadrupoles for transverse beam focusing, the spiral inflector to bring the beam from the vertical axis onto the median plane, and the specific RF electrode (aka “dees”) shapes to center the beam during the first turns.
  • The main magnet and RF Cavities for acceleration to 60 MeV/amu.
  • (not shown) The extraction system consisting of electrostatic and magnetic channels guiding the beam out of the cyclotron with the goal of extracting 5 mA of $\mathrm{H}_{2}^{+}$.